The length of a certain rectangle is 4 m greater than five times its width. Find the dimensions of the rectangle if the perimeter is 20 m. (Hint: the perimeter of a rectangle is P = 2L + 2W.)
Found 2 solutions by rfer, ikleyn:
Answer by rfer(16322)

2(5w+4)+2w=20
10w+8+2w=20
12w=20-8
w=12/12
w=1
l=5*1+4
l=9

Answer by ikleyn(52779)